# MambaOut: Do We Really Need Mamba for Vision?

### v2

@@ 1 Introduction
In this paper, we investigate the nature of Mamba, and conceptually summarizes that Mamba is ideally suited for tasks with two key characteristics: long-sequence and autoregressive, because of the the inherent RNN mechanism of SSM [27, 26, 25] (see explanation of Figure 2 and Figure 3). Unfortunately, not many vision tasks possess both characteristics. Image classification on ImageNet, for example, conforms to neither, while object detection & instance segmentation on COCO and semantic segmentation on ADE20K conform to only long-sequence. Autoregressive characteristic, on the other hand, demands that each token aggregate information solely from preceding and current tokens, a concept denoted as causal mode for token mixing.[63] (see Figure 3(a)). In fact, all visual recognition tasks fall within the understanding domain rather than the generative one, meaning that the model can see the entire image at once. As such, imposing additional causal constraints on token mixing in visual recognition models could lead to a performance drop (see Figure 3(b)). Although this issue can be mitigated via bidirectional branches [68], it is inevitable that the issue persists within each branch.
@@ 2 Related work
Transformer has been widely utilized in various domains, like BERT [20] and GPT series [60, 61, 6, 1] in NLP and ViT [23] in computer vision. However, the attention module in Transformers scales quadratically with sequence length, presenting a significant computational challenge. Numerous studies [72] have explored various strategies to mitigate this issue, including low-rank approaches [78], kernelization [40, 12], token mixing range limitation [5, 97, 51, 29], and history memory compression [62]. More recently, RNN-like methods [17, 40, 98], particularly RWKV [59] and Mamba [25], have garnered attention for their promising results in large language models [59, 47].
@@
Unlike these initiatives, our work does not aim to design new visual Mamba models. Instead, we explore a pertinent research question about the necessity of Mamba [25] in traditional visual recognition contexts [19, 66, 48, 103]. We hope this paper can provide insights for future research on visual Mamba models.
@@ 3.4 Hypotheses regarding the necessity of Mamba for vision
  • Hypothesis 2: It is still worthwhile to further explore the potential of SSM for visual detection and segmentation since these tasks align with Characteristic 2, despite not fulfilling Characteristic 1.
@@ 4.1 Gated CNN and MambaOut
Next, we aim to validate our hypotheses empirically. As depicted in Figure 1(a), Mamba block is based [25] on the Gated CNN block [18]. Both the Gated CNN and Mamba’s meta-architecture can be considered as a simplified integration of the MetaFormer’s [91] token mixer and an MLP, akin to MetaNeXt [93]. Formally, given the input $X\in\mathbb{R}^{N\times D}$, the meta-architecture is formulated as:
@@
where $\mathrm{Norm}(\cdot)$ represents normalization [38, 2, 83]; $\mathrm{TokenMixer}(\cdot)$ $\mathrm{TokenMixer}(\cdot)$ refers to the module to conduct token mixing [92]; $W_{1}\in\mathbb{R}^{D\times rD}$, $W_{2}\in\mathbb{R}^{D\times rD}$ and $W_{3}\in\mathbb{R}^{rD\times D}$ are learnable parameters with MLP expansion $r$; $\sigma$ is activation function [24, 33]. Token mixers of Gated CNN and Mamba are:
@@ 4.2 Image classification on ImageNet
Setup. ImageNet [19, 66] serves as the gold standard benchmark for image classification, encompassing a wide array of 1,000 common classes. It comprises approximately 1.3 million training images and 50,000 validation images. The training scheme follows DeiT [73] without distillation. Specifically, the used data augmentation contains random resized crop (input image size of $224^{2}$), horizontal flip, RandAugment [15], Mixup [100], CutMix [96], Random Erasing [102] and color jitter, and regularization techniques include weight decay, stochastic depth [36] and label smoothing [71]. All our models are trained by AdamW [53, 41]. The learning rate scaling rule is $\mathrm{lr}=\frac{\mathrm{batchsize}}{1024}\times 10^{-3}$. In this paper, we set the batch size to 4096, so the learning rate is $0.004$. Our MambaOut models are implemented with PyTorch [56] and timm [81] libraries and trained on TPU v3. More training hyper-parameters are shown in Table 5 in the appendix.
Results. The performance of our MambaOut models, visual Mamba models, and various other convolution and attention-based models on ImageNet [19, 66] is presented in Table 1. Notably, our MambaOut models, which do not incorporate SSM, consistently outperform visual Mamba models [104, 50, 37, 58, 88] that include SSM across all model sizes. For instance, the MambaOut-Small model achieves top-1 accuracy of 84.1%, 0.4% higher than that of LocalVMamba-S [37], while requiring only 79% of the MACs. These results strongly support our Hypothesis 1, which posits that introducing SSM for image classification on ImageNet is unnecessary, aligning with the principle of Occam’s razor.
Additionally, visual Mamba models currently exhibit a significant performance gap when compared to state-of-the-art convolution and attention models. For instance, the CAFormer-M36 [92], which employs traditional token mixers like simple separable convolutions [67] and standard attention mechanisms [76], outperforms all visual Mamba models of comparable size by more than 1% in accuracy. Should future research aim to challenge our Hypothesis 1, it will be necessary to develop visual Mamba models with token mixers of convolution and SSM to achieve state-of-the-art performance on ImageNet.
@@ 4.3 Object detection & instance segmentation on COCO
TransNeXt-Tiny [69] Conv + Attn 48 49.9 71.5 54.9 44.6 68.6 48.1
@@
TransNeXt-Small [69] Conv + Attn 69 51.1 72.6 56.2 45.5 69.8 49.1
@@
TransNeXt-Base [69] Conv + Attn 109 51.7 73.2 56.9 45.9 70.5 49.7
@@
TransNeXt-Tiny [69] Conv + Attn 59 51.1 51.7
@@
TransNeXt-Small [69] Conv + Attn 80 52.2 52.8
@@
TransNeXt-Base [69] Conv + Attn 90 53.0 53.7
@@ 5 Conclusion
In this paper, we discuss the Mamba mechanism conceptually and conclude that it is ideally suited for tasks with long-sequence and autoregressive characteristics. We analyze common visual tasks against these criteria and argue that introducing Mamba for ImageNet image classification is unnecessary, as it meets neither characteristic. However, the potential of Mamba for visual detection and segmentation tasks, which align with the long-sequence characteristic, merits further exploration. To substantiate our claims empirically, we develop MambaOut models that employ Mamba blocks without their core token mixer, SSM. MambaOut surpasses all visual Mamba models on ImageNet, yet it exhibits a notable performance gap compared to state-of-the-art visual Mamba models, thereby validating our assertions. Due to computational resource limitations, this paper only verifies the Mamba concept for visual tasks. In the future, we may further explore Mamba and RNN concepts as well as the integration of RNN and Transformers for large language models (LLMs) and large multimodal models (LMMs).

### v3

@@ 1 Introduction
In this paper, we investigate the nature of Mamba, and conceptually summarize that Mamba is ideally suited for tasks with two key characteristics: long-sequence and autoregressive, because of the inherent RNN mechanism of SSM [27, 26, 25] (see explanation of Figure 2 and Figure 3). Unfortunately, not many vision tasks possess both characteristics. Image classification on ImageNet, for example, conforms to neither, while object detection & instance segmentation on COCO and semantic segmentation on ADE20K conform only to the long-sequence. Autoregressive characteristic, on the other hand, demands that each token aggregate information solely from preceding and current tokens, a concept denoted as causal mode for token mixing [63] (see Figure 3(a)). In fact, all visual recognition tasks fall within the understanding domain rather than the generative one, meaning that the model can see the entire image at once. As such, imposing additional causal constraints on token mixing in visual recognition models could lead to a performance drop (see Figure 3(b)). Although this issue can be mitigated via bidirectional branches [68], it is inevitable that the issue persists within each branch.
@@ 2 Related work
Transformer has been widely utilized across various domains, including BERT [20] and GPT series [60, 61, 6, 1] in NLP and ViT [23] in computer vision. However, the attention module in Transformers scales quadratically with sequence length, presenting a significant computational challenge. Numerous studies [72] have explored various strategies to mitigate this issue, including low-rank approaches [78], kernelization [40, 12], token mixing range limitation [5, 97, 51, 29], and history memory compression [62]. More recently, RNN-like methods [17, 40, 98], particularly RWKV [59] and Mamba [25], have garnered attention for their promising results in large language models [59, 47].
@@
Unlike these initiatives, our work does not aim to design new visual Mamba models. Instead, we explore a pertinent research question about the necessity of Mamba [25] in visual recognition contexts [19, 66, 48, 103]. We hope this paper can provide insights for future research on visual Mamba models.
@@ 3.4 Hypotheses regarding the necessity of Mamba for vision
  • Hypothesis 2: It is still worthwhile to further explore the potential of SSM for visual detection and segmentation since these tasks align with Characteristic 1, despite not fulfilling Characteristic 2.
@@ 4.1 Gated CNN and MambaOut
Next, we aim to validate our hypotheses empirically. As depicted in Figure 1(a), Mamba block [25] is based on the Gated CNN block [18]. The meta-architecture of Gated CNN and Mamba can be considered as a simplified integration of the MetaFormer’s [91] token mixer and an MLP, akin to MetaNeXt [93]. Formally, given the input $X\in\mathbb{R}^{N\times D}$, the meta-architecture is formulated as:
@@
where $\mathrm{Norm}(\cdot)$ represents normalization [38, 2, 83]; $\mathrm{TokenMixer}(\cdot)$ refers to the module to conduct token mixing [92]; $W_{1}\in\mathbb{R}^{D\times rD}$, $W_{2}\in\mathbb{R}^{D\times rD}$ and $W_{3}\in\mathbb{R}^{rD\times D}$ are learnable parameters with MLP expansion $r$; $\sigma$ is activation function [24, 33]. Token mixers of Gated CNN and Mamba are:
@@ 4.2 Image classification on ImageNet
Setup. ImageNet [19, 66] serves as the gold standard benchmark for image classification, encompassing a wide array of 1,000 common classes. It comprises approximately 1.3 million training images and 50,000 validation images. The training scheme follows DeiT [73] without distillation. Specifically, the used data augmentation contains random resized crop (input image size of $224^{2}$), horizontal flip, RandAugment [15], Mixup [100], CutMix [96], Random Erasing [102] and color jitter; Regularization techniques include weight decay, stochastic depth [36] and label smoothing [71]. All our models are trained by AdamW [53, 41]. The learning rate scaling rule is $\mathrm{lr}=\frac{\mathrm{batchsize}}{1024}\times 10^{-3}$. In this paper, we set the batch size to 4096, so the learning rate is $0.004$. Our MambaOut models are implemented with PyTorch [56] and timm [81] libraries and trained on TPU v3. More training hyper-parameters are shown in Table 5 in the appendix.
Results. The performance of our MambaOut models, visual Mamba models, and other various convolution and attention-based models on ImageNet [19, 66] is presented in Table 1. Notably, our MambaOut models, which do not incorporate SSM, consistently outperform visual Mamba models [104, 50, 37, 58, 88] that include SSM across all model sizes. For instance, the MambaOut-Small model achieves top-1 accuracy of 84.1%, 0.4% higher than that of LocalVMamba-S [37], while requiring only 79% of the MACs. These results strongly support our Hypothesis 1, which posits that introducing SSM for image classification on ImageNet is unnecessary, aligning with the principle of Occam’s razor.
Additionally, visual Mamba models currently exhibit a significant performance gap when compared to state-of-the-art convolution and attention models. For instance, the CAFormer-M36 [92], which employs old-fashioned token mixers of simple separable convolutions from MobileNetV2 [67] and vanilla attention from Transformer [76] invented more than 7 years ago, outperforms all visual Mamba models of comparable size by more than 1% accuracy. Should future research aim to challenge our Hypothesis 1, it will be necessary to develop visual Mamba models with token mixers of convolution and SSM to achieve state-of-the-art performance on ImageNet.
@@ 4.3 Object detection & instance segmentation on COCO
TransNeXt-Tiny [69] Conv + Attn 48 356 49.9 71.5 54.9 44.6 68.6 48.1
@@
TransNeXt-Small [69] Conv + Attn 69 516 51.1 72.6 56.2 45.5 69.8 49.1
@@
TransNeXt-Base [69] Conv + Attn 109 728 51.7 73.2 56.9 45.9 70.5 49.7
@@
TransNeXt-Tiny [69] Conv + Attn 59 978 51.1 51.2
@@
TransNeXt-Small [69] Conv + Attn 80 1089 52.2 52.3
@@
TransNeXt-Base [69] Conv + Attn 121 1268 53.0 53.4
@@ 5 Conclusion
In this paper, we discuss the Mamba mechanism conceptually and conclude that it is ideally suited for tasks with long-sequence and autoregressive characteristics. We analyze common visual tasks against these criteria and argue that introducing Mamba for ImageNet image classification is unnecessary, as it meets neither characteristic. However, the potential of Mamba for visual detection and segmentation tasks, which align with at least the long-sequence characteristic, merits further exploration. To substantiate our claims empirically, we develop MambaOut models that employ Mamba blocks without their core token mixer, SSM. MambaOut surpasses all visual Mamba models on ImageNet, yet it exhibits a notable performance gap compared to state-of-the-art visual Mamba models, thereby validating our assertions. Due to computational resource limitations, this paper only verifies the Mamba concept for visual tasks. In the future, we may further explore Mamba and RNN concepts as well as the integration of RNN and Transformer for large language models (LLMs) and large multimodal models (LMMs).